The birth weight of babies is normally distributed with a mean of 3.8 kg. A researcher suspects that the weight of babies from mothers who smoked a lot during pregnancy will be lower than the population mean. To examine this, he takes a random sample of 26 babies from mothers who smoked a lot. The mean birthweight in this group was calculated as 3.48 kg with an estimated standard deviation of 0.80 kg. He takes Ho: mu = 3.8 kg, H1: mu < 3.8 kg. According to this data, what is the p-value for rejecting Ho? If alpha = 5% is used, will Ho be rejected?
Answer #1

Solution :

This is the left tailed test .

The null and alternative hypothesis is ,

H0 :   = 3.8

Ha : < 3.8

= 3.48

= 3.8

s = 0.80

n = 26

degrees of freedom = n - 1 = 26 - 1 = 25

Test statistic = t

= ( - ) / s / n

= (3.48 - 3.8) / 0.80 / 26

= -2.04

Test statistic = -2.04

P-value = 0.026

= 0.05

P-value <

Reject the null hypothesis .

H0 is rejected .